OpenStudy (anonymous):

Find a Cartesian equation for the curvegiven in parametric form byx(t) = 4 cos2t , y(t) = 3 sin2t .

OpenStudy (experimentx):

I think eliminating t from those two equations will do.

OpenStudy (anonymous):

no more complicated...@ experimentx...

OpenStudy (dumbcow):

well the straightforward way is solve x(t) for 2t and substitute it in for 2t in y(t) \[2t = \cos^{-1} (x/4)\] \[y = 3\sin (\cos^{-1} (x/4)) = 3\sqrt{1-\cos^{2} (\cos^{-1} (x/4))}\] \[y = 3\sqrt{1-\frac{x^{2}}{16}}\]

OpenStudy (experimentx):

x = 4 cos2t x/4 = cos2t (x/4)^2 = (cos2t)^2 ---- 1 y = 3 sin2t y/3 = sin2t (y/3)^2 = (sin2t)^2 -----2 add 1 and 2 (y/3)^2+(x/4)^2 = (sin2t)^2 + (cos2t)^2 = 1 (y/3)^2+(x/4)^2 = 1 An ellipse.

OpenStudy (anonymous):

well according to the answer choices 1.. 4x + 3 y = 12 2.x/4 -y/3=1/12 . 3x − 4 y = 12

OpenStudy (anonymous):

lemme type the other ones

OpenStudy (dumbcow):

@experimentX, thats the better or more efficient way of doing it :)

OpenStudy (anonymous):

4.x/3-y/4=1/12 5.3x + 4 y = 12 6.x/3+4y=12

OpenStudy (experimentx):

thanks ... learned on OS. and your's is my classic and always working one.

OpenStudy (dumbcow):

torken, none of the choices have x or y squared ??

OpenStudy (anonymous):

nope

OpenStudy (dumbcow):

its def not a linear equation

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

x(t)=4cos^2 2t y(t)=3sin^2 2t

OpenStudy (anonymous):

i forgot to put the cos^2 and sin ^2 >.<

OpenStudy (experimentx):

bad luck ... just add them, you will have lines.

OpenStudy (dumbcow):

haha

OpenStudy (dumbcow):

x/4 + y/3 = 1 3x +4y = 12

OpenStudy (anonymous):

thx much hehe sooo confusing questions ><